# literature.bib

### v1

@@ 6 Conclusion and Future Work
We conducted detailed performance studies to understand the performance characteristics of the hybrid AD approach both with respect to memory usage and AD-specific runtime. We identified the strategy for identifier management, preaccumulations, and atomic updates on adjoint variables as key differences to the previous MPI-parallel differentiation strategy in SU2 and quantified their impact on the performance in various test cases. Throughout all tests, we observed that OpenMP consistently leads to better memory scaling. As we compared the hybrid AD approach to a non-hybrid AD approach with equivalent identifier management and preaccumulation, we could demonstrate a significant reduction in memory usage for highly parallel setups, which is an important insight given the substantial memory cost of tape-based reverse AD. However, we also observed notable runtime overheads of hybrid AD, especially compared to an MPI approach with linear management and all preaccumulations, so that hybrid AD could only serve as a tool to trade memory for runtime in our test cases. With the shared reading optimization and optimized adjoint vector management, we already applied two initial improvements of hybrid AD and showcased their effectiveness in our performance tests. However, these only mark the first steps towards efficient hybrid parallel discrete adjoints in SU2. Right now, hybrid parallel discrete adjoints could be useful for problems that exceed available memory capacities, especially on CPUs with large NUMA domains and high degrees of OpenMP parallelism, and for large numbers of discrete adjoint iterations.
@@
Our observations suggest various perspectives for future improvement. Within SU2, it could be worthwhile to investigate and, if possible, improve the scalability of the recordings, even though it is not the dominant part of the AD-specific runtime and does not increase with the number of discrete adjoint iterations. The evaluation time and its scalability could be improved by eliminating more atomic updates, either by identifying further exclusive read access or by revising the data access patterns to reduce non-exclusive read access. Although the scaling factors for evaluations are competitive already, they would need to be larger (or break down later) than the MPI scaling factors so that hybrid AD outperforms MPI-parallel AD not only in terms of memory but also in terms of runtime for high degrees of parallelism. Our studies suggest that the parallel reuse identifier management that is currently in place for the hybrid AD approach might not be the best choice for SU2. It could be beneficial to move to a parallel single-use scheme that supports optimized copies, with the additional benefit that it would readily support the multizone taping approach, but with the drawback that copies behave like reference in terms of exclusive read access. We also identified preaccumulation as a promising target for reductions of the memory usage. Re-enabling it in the presence of non-exclusive read access, however, is likely accompanied by new memory and/or runtime overheads in the recording phase. We plan to evaluate these improvements in follow-up studies.
@@ Acknowledgements
Johannes Blühdorn, Max Aehle, and Nicolas R. Gauger gratefully acknowledge the funding of the German National High Performance Computing (NHR) association for the Center NHR South-West.
[…]

### v2

@@ 5 Conclusion and future work
We conducted detailed performance studies to understand the performance characteristics of the hybrid AD approach both with respect to memory usage and AD-specific runtime. We identified the strategy for identifier management, preaccumulations, and atomic updates on adjoint variables as key differences to the previous MPI-parallel differentiation strategy in SU2 and quantified their impact on the performance in various test cases. Throughout all tests, we observed that OpenMP consistently leads to better memory scaling. As we compared the hybrid AD approach to a non-hybrid AD approach with equivalent identifier management and preaccumulation, we could demonstrate a significant reduction in memory usage for highly parallel setups, which is an important insight given the substantial memory cost of tape-based reverse AD. However, we also observed notable runtime overheads of hybrid AD, especially compared to an MPI approach with linear management and all preaccumulations, so that hybrid AD could only serve as a tool to trade memory for runtime in our test cases. With the shared reading optimization and optimized adjoint vector management, we already applied two initial improvements of hybrid AD and showcased their effectiveness in our performance tests. However, these only mark the first steps towards efficient hybrid parallel discrete adjoints in SU2. Right now, hybrid parallel discrete adjoints could be useful for problems that exceed available memory capacities, especially on CPUs where large numbers of OpenMP threads per MPI process can be used efficiently, and for large numbers of discrete adjoint iterations.
@@
Our observations suggest various perspectives for future improvement. Within SU2, it could be worthwhile to investigate and, if possible, improve the scalability of the recordings, even though it is not the dominant part of the AD-specific runtime and does not increase with the number of discrete adjoint iterations. The evaluation time and its scalability could be improved by eliminating more atomic updates, either by identifying further exclusive read access or by revising the data access patterns to reduce non-exclusive read access. Although the scaling factors for evaluations are competitive already, they would need to be larger (or break down later) than the MPI scaling factors so that hybrid AD outperforms MPI-parallel AD not only in terms of memory but also in terms of runtime for high degrees of parallelism. Our studies suggest that the parallel reuse identifier management that is currently in place for the hybrid AD approach might not be the best choice for SU2. It could be beneficial to move to a parallel single-use scheme that supports optimized copies, with the additional benefit that it would readily support the multizone taping approach, but with the drawback that copies behave like reference in terms of exclusive read access. We plan to evaluate these improvements in further studies. As we disabled simultaneous preaccumulations with shared inputs, we expect that approaches that re-enable them lead to reduced memory usage, however, likely accompanied by new memory and/or runtime overheads in the recording phase. This issue is addressed in detail in follow-up research in [38], where we propose and evaluate approaches that re-enable these preaccumulations.
@@ Acknowledgements
Johannes Blühdorn, Max Aehle, and Nicolas R. Gauger gratefully acknowledge funding from the German National High Performance Computing (NHR) association for the Center NHR South-West.